OpenStudy (anonymous):

How do I determine the amplitude and period of this? y= sin 4x

OpenStudy (anonymous):

the amplitude is the numerical coefficient that precedes the sine function. In this case, it is 1 since there is no number preceding the sin. The period is 2pi/the numerical coefficient of x. In this case it is 4

OpenStudy (zepp):

The period of a sin function could be found by using the formula \[\frac{b}{2 \pi}\]

OpenStudy (anonymous):

amplitude = maximum - equation of axis

OpenStudy (anonymous):

y=asin(kx) our a is the amplitude in this case our a is 1 so our amplitude is 1

OpenStudy (anonymous):

our period = 360/k our k in this case is 4 360/4=90 degrees so our period is 90 degrees

OpenStudy (zepp):

And the amplitude would be 4. Let's take a look at the basic sin function, \[f(x) = sin(x)\]  We know that the max and the min of this function is -1 and 1, amplitude 1 Then the constant changed for 4 \[4*1 =4\]